OpenStudy (anonymous):

divide: 8xy^2-14y^2/2y explain.

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

ok

OpenStudy (anonymous):

so do i put the common numbers together first?

OpenStudy (anonymous):

like hold on let me try it

OpenStudy (danjs):

\[\frac{ 8*x*y^2 - 14*y }{ 2y }\]

OpenStudy (danjs):

recall: \[\frac{ 1 }{ x^a } = \frac{ x ^{-a} }{ 1 }\]

OpenStudy (mertsj):

\[\frac{8xy^2-14y^2}{2y}=\frac{8xy^2}{2y}-\frac{14y^2}{2y}\]

OpenStudy (anonymous):

-6xy^2/2y= -3xy????

OpenStudy (anonymous):

no lol i need to work for it or ill never learn

OpenStudy (mertsj):

Now reduce each fraction

OpenStudy (anonymous):

so simplify each fraction???

OpenStudy (mertsj):

yes

OpenStudy (danjs):

\[\frac{ y^2 }{ y } = y^2*y^{-1} = y ^{2-1} = y\]

OpenStudy (danjs):

you make the exponent negative when you move it to the other side of the fraction

OpenStudy (anonymous):

wait do i reduce 8/2 and 14/2????

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

oh okay hold on

OpenStudy (anonymous):

so that means i see what all can go into the numerator and denominator right??

OpenStudy (mertsj):

What is 8 divided by 2?

OpenStudy (mertsj):

What is 14 divided by 2?

OpenStudy (anonymous):

oh its 7 and 16 and so then i divide it and i get a decimal

OpenStudy (anonymous):

well the choices are A.4xy^2-7y^2 S.4x-7 C.4xy-7y D.4y-7x

OpenStudy (anonymous):

but my answer ended up as 7/16 or a big decimal

OpenStudy (danjs):

4xy - 7y

OpenStudy (danjs):

here.....

OpenStudy (danjs):

\[\frac{ 8xy^2 - 14y^2}{ 2y } = \frac{ 8xy^2 }{ 2y } + \frac{ 14y^2 }{ 2y } = \frac{ 4xy }{ 1 }+ \frac{ 7y }{ 1 }\]

OpenStudy (danjs):

so just... 4xy + 7y

OpenStudy (danjs):

oh its minus

OpenStudy (danjs):

replace those + with -

OpenStudy (danjs):

4xy - 7y

OpenStudy (anonymous):

oh okay so when theres a positive sign i change it to a negative?

OpenStudy (danjs):

no the original problem was minus, i just misstyped it

OpenStudy (anonymous):

oh okay ill tag you again

OpenStudy (danjs):

alright